A simple harmonic one dimensional oscillator has energy level is given by En=(n+{1÷2}h cross omega.where omega is angular frequency and h=0,1,2,3,.... .Suppose that this oscillator is in thermal contact with a heat reservior at tempararature T low enough so that kT<<hcross omega.Find the ratio of probability of oscillator being in first excited state to the probability of its being in the ground state.
1
Expert's answer
2020-03-30T07:58:07-0400

As per the given question,

Energy of the one dimensional simple harmonic oscillator "E_n=(n+\dfrac{1}{2})\hbar\omega"

Now, we know that probability of the oscillator in nth excited state,

"P=e^{-\beta E_n}"

Now,

probability of the oscillator being in the first excited state "P_1=e^{-\beta E_1}"

probability of the oscillator being in the ground state "P_0=e^{-\beta E_0}"

Hence, the required ratio "\dfrac{P_1}{P_o}=\dfrac{e^{-\beta E_1}}{e^{-\beta E_0}}=\dfrac{e^{-\beta (1+\frac{1}{2})\hbar\omega}}{e^{-\beta (0+\frac{1}{2})\hbar\omega}}"

"=\dfrac{e^{-\beta (\frac{3}{2})\hbar\omega}}{e^{-\beta (\frac{1}{2})\hbar\omega}}"


"=e^{-\beta\hbar\omega(-\frac{3}{2}+\frac{1}{2})}"


"=e^{-\beta \hbar\omega}"